9. Find I? 10 v 1도 210 0 3100 2SO | 110V 210 0 | 3 3100
s=25
Answer #1

we know source transformation technique use this technique for this Given Circuit 10V apply so Transformation {252 s 202 Ilio

TOXIUS connected in parallel connection so Combine Totio 3250 LOL . loxlo A css 10+10 Connected in parallel so combine - tor

LoD . 1852 = = 1X5 =SV § 352 Rz lor + lov 35+35+5=451 +(5+10-5) = slov lor By again sowoce Transfromation

45.410. (452||10-2) = 8.182 102 45+10 45+10 0.224 8.182 v slor V -0.22A By current Division Rule .8.18 I = -0.22% 8:18+10 I =